### v3

# CVBench: Benchmarking Cross-Video Synergies for Complex Multimodal Reasoning

Abstract
While multimodal large language models (MLLMs) exhibit strong performance on single-video tasks (e.g., video question answering), their capability for spatiotemporal pattern reasoning across multiple videos remains a critical gap in pattern recognition research. However, this capability is essential for real-world applications, including multi-camera surveillance and cross-video procedural learning. To bridge this gap, we present CVBench, the first diagnostic benchmark designed to assess cross-video relational reasoning rigorously. CVBench comprises 1,000 question-answer pairs spanning three hierarchical tiers: cross-video object association (identifying shared entities), cross-video event association (linking temporal or causal event chains), and cross-video complex reasoning (integrating commonsense and domain knowledge). Built from five domain-diverse video clusters (e.g., sports, life records), the benchmark challenges models to analyze and integrate spatiotemporal patterns from dynamic visual streams. Extensive evaluation of 10+ leading MLLMs (including GPT-4o, Gemini-2.0-flash, Qwen2.5-VL) under zero-shot or chain-of-thought prompting paradigms. Key findings reveal stark performance gaps: even top models, such as GPT-4o, achieve only 63.5% accuracy on causal reasoning tasks, compared to the 91.3% accuracy of human performance. Crucially, our analysis reveals fundamental bottlenecks inherent in current MLLMs architectures, notably deficient inter-video context retention and poor disambiguation of overlapping entities. CVBench establishes a rigorous framework for advancing pattern recognition methodologies in multi-video scenarios, providing architectural insights for next-generation models. The data and evaluation code are available at: https://github.com/Hokhim2/CVBench.
Abstract
While multimodal large language models (MLLMs) exhibit strong performance on single-video tasks (e.g., video question answering), their ability across multiple videos remains critically underexplored. However, this capability is essential for real-world applications, including multi-camera surveillance and cross-video procedural learning. To bridge this gap, we present CVBench, the first comprehensive benchmark designed to assess cross-video relational reasoning rigorously. CVBench comprises 1,000 question-answer pairs spanning three hierarchical tiers: cross-video object association (identifying shared entities), cross-video event association (linking temporal or causal event chains), and cross-video complex reasoning (integrating commonsense and domain knowledge). Built from five domain-diverse video clusters (e.g., sports, life records), the benchmark challenges models to synthesise information across dynamic visual contexts. Extensive evaluation of 10+ leading MLLMs (including GPT-4o, Gemini-2.0-flash, Qwen2.5-VL) under zero-shot or chain-of-thought prompting paradigms. Key findings reveal stark performance gaps: even top models, such as GPT-4o, achieve only 60% accuracy on causal reasoning tasks, compared to the 91% accuracy of human performance. Crucially, our analysis reveals fundamental bottlenecks inherent in current MLLM architectures, notably deficient inter-video context retention and poor disambiguation of overlapping entities. CVBench establishes a rigorous framework for diagnosing and advancing multi-video reasoning, offering architectural insights for next-generation MLLMs.The data and evaluation code are available at https://github.com/Hokhim2/CVBench.
keywords:
Multimodal large language models, Cross-video reasoning, Video question answering, Multivideo understanding, Benchmark evaluation.
journal: Pattern Recognition
[…]
[label1]organization=School of Systems Science and Engineering, Sun Yat-sen University,addressline=No. 135 Xingang West Road, city=Guangzhou, postcode=510275, state=Guangdong, country=China
[…]
[label2]organization=Tencent,addressline=Keji Zhongyi Road, Nanshan District, city=Shenzhen, postcode=518057, state=Guangdong, country=China
[…]
[label3]organization=School of Electronics and Communication Engineering, Sun Yat-sen University,addressline=No. 66 Gongchang Road, city=Shenzhen, postcode=518107, state=Guangdong, country=China
[…]
[label4]organization=School of Electronics and Information Engineering, Foshan University,addressline=No. 18 Jiangwan 1st Road, city=Foshan, postcode=528225, state=Guangdong, country=China
[…]
[label5]organization=Department of Computer Science and Engineering, Southern University of Science and Technology, addressline=1088 Xueyuan Avenue, city=Shenzhen, postcode=518055, state=Guangdong, country=China
{highlights}
First systematic benchmark for evaluating spatiotemporal relational pattern recognition across multiple videos, featuring a hierarchical structure with 1,000 expert-annotated QA pairs to diagnose capabilities in object association, event association, and complex reasoning.
Comprehensive evaluation reveals significant limitations of current models: even state-of-the-art multimodal models achieve only 63.5% accuracy on causal reasoning, far below human performance of 91.3%, exposing fundamental bottlenecks in inter-video pattern integration and entity disambiguation.
Provides actionable architectural insights for advancing pattern recognition models, including cross-video memory mechanisms and grounded spatiotemporal reasoning, offering a clear pathway toward robust multi-video understanding systems.
1 Introduction
The ability to reason across multiple video streams is fundamental to real-world applications, such as multi-camera surveillance systems, distributed procedural learning, and multi-view activity analysis Wu2025 . While multimodal large language models (MLLMs) have demonstrated significant progress in single-video understanding CAI2025111670 , their capacity for spatiotemporal pattern reasoning across videos remains a critical and underexplored challenge in pattern recognition research ZHENG2025111319 . This capability requires models to identify and analyze relational patterns across distinct spatiotemporal contexts 11008449 , such as entity correspondences, event sequences, and causal links. The research gap stems not merely from a lack of single-video benchmarks, but from the inherent limitations of existing evaluations in capturing cross-video relational dynamics, which are central to real-world pattern recognition systems.
Current MLLMs face fundamental barriers in cross-video scenarios: 1) Inter-video context linking: models lack mechanisms to persistently integrate and retain object/event states cross-video separated by time and space Li10658165 , 11093290 ; 2) Entity disambiguation: overlapping or visually similar entities cross-video lack unique spatiotemporal grounding, causing identification errors; 3) Temporal-causal modeling: existing architectures poorly capture long-range dependencies and causal links spanning asynchronous video sequences 11095110 .
Figure 1: Illustration of CVBench task definition. CVBench focuses on three core tasks that are crucial for real-world applications: a) cross-video object association (identifying shared entities across different videos), b) cross-video event association (linking temporal or causal event chains that span across video sequences), and c) cross-video complex reasoning (integrating commonsense to understand and reason about complex interrelationships between video content). The benchmark format is multi-choice QA, and here we only show the correct answer for visualization.
To address these limitations and provide a diagnostic framework for the community, we introduce CVBench, the first comprehensive benchmark designed specifically for evaluating cross-video relational pattern recognition. As illustrated in Fig. 1, CVBench structures the evaluation around three hierarchically organized tasks:cross-video object association for verifying entity persistence across varying visual contexts, cross-video event association for modeling temporal sequences and causal relationships cross-video, and cross-video complex reasoning for integrating visual cues and external knowledge for hierarchical spatiotemporal understanding. CVBench comprises 1,000 QA pairs sampled from five diverse domains (artistic performances, sports competitions, films and television, life records, and knowledge) to ensure generalization. This domain diversity challenges models to reason across dynamic contexts, requiring spatio-temporal integration and multi-view synthesis. Expert annotations capture real-world complexities by modeling entities, events, and inter-video relationships. This rigorous foundation enables diagnostic evaluation of cross-video reasoning capabilities.
We evaluate over ten state-of-the-art closed-sourced and open-sourced MLLMs, including GPT-4o, Gemini series, and Qwen2.5-VL, under zero-shot and chain-of-thought prompting paradigms. Our evaluation reveals significant performance gaps in these models. Specifically, GPT-4o achieves only 60.2% accuracy on causal reasoning tasks, which is lower than the human-level performance of 91.3%. Additionally, these models struggle with inter-video context retention and difficulties in disambiguating overlapping entities, which hinders their ability to integrate information from multiple sources. These findings underscore the limitations of current MLLMs in handling complex cross-video tasks, highlighting the need for more advanced architectures to address challenges like long-term context retention and entity disambiguation.
The contributions of this paper can be summarized as:
1) We introduce CVBench, the first diagnostic benchmark explicitly targeting cross-video understanding through object association, event association, and complex reasoning.
2) We release a carefully curated multi-domain dataset with comprehensive annotations, enabling rigorous stress-testing of spatiotemporal integration and multi-view synthesis capabilities.
3) We provide a systematic analysis of state-of-the-art MLLMs on CVBench, exposing critical performance gaps and identifying key research directions for advancing cross-video pattern reasoning in the field.
2 Related work
Multimodal large language models
The evolution of MLLMs has progressed from image-based Zhang10547418 to video-capable architectures Fu11093290 , ZHANG2026111925 , marked by two main improvements: 1) Long-range context modeling that enables efficient handling excessively long temporal contexts Yue11093104 , Song10657734 ; 2) Spatio-temporal fusion that integrates both appearance and motion dynamics with language reasoning Ren10656135 , Zhang11094660 , Liu11094049 . Despite these advances, current MLLMs remain architecturally constrained to single-video scenario, lacking mechanisms for cross-video interactions, exhibiting inadequate spatiotemporal grounding for overlapping entities, and demonstrating limited modeling of temporal event chains. These fundamental limitations restrict their applicability to multi-camera surveillance and cross-video applications Yang11037495 .
Video understanding benchmarks
A comprehensive and objective benchmark is essential for evaluating MLLMs, enabling the comparison and investigation of the performance of various models SHAO2026112724 . Recent advances in multimodal learning have led to the development of several benchmarks for video understanding tasks ZHONG2025111035 . However, existing video evaluation frameworks exhibit critical limitations in assessing cross-video reasoning capabilities. First, general multimodal frameworks are predominantly with broad domain coverage, such as Video-MMMU and Video-MME Fu11093290 ), to evaluate instruction-following capabilities but lacking cross-video relational tasks. Second, long-term video benchmarks contain longer videos with multiple sub-events, however, are confined to single source or similar creators, such as VideoTree Wang11094563 , Video-Bench Ning11284911 , and Chapter-Llama Ventura11092338 . These deficiencies fundamentally limit their applicability to real-world scenarios in multi-camera surveillance and cross-video procedural learning.
To address these gaps, we introduce CVBench, the first diagnostic framework for cross-video relational reasoning across different sources. CVBench pioneers a three-tiered evaluation of spatiotemporal integration capabilities, encompassing object-level association, event-level association and complex reasoning.
3 CVBench: cross-video reasoning benchmark
3.1 Task definition
CVBench pioneers the evaluation of cross-video relational reasoning through a three-tiered diagnostic framework, encompassing object association, event association, and complex reasoning. This taxonomy systematically addresses critical gaps in real-world video understanding by requiring models to establish inter-video relationships through spatiotemporal pattern recognition beyond single-video analysis. The framework comprises 15 subtasks (detailed in Appendix) categorized as follows:
1) Object association: Cross-video object recognition, multi-video attribute recognition, joint-video counting, and cross-video entity matching. These require consistent identification of entities across domains or viewpoints despite occlusion or deformation.
2) Event association: Cross-video anomaly detection, scene recognition, key-action recognition, event retrieval, multi-view scene understanding, and temporal reasoning. These demand fusion of spatiotemporal signatures across videos to reconstruct event dynamics.
3) Complex reasoning: Joint-video spatial navigating, video difference captioning, cross-video counterfactual reasoning, joint-video summarization, and procedural transfer. These necessitate higher-order synthesis including spatial imagination, fine-grained reasoning, causal inference and adaptive knowledge transfer.
3.2 Dataset collection
CVBench is constructed with three-step process: video curation, question-answer annotation, and quality control. First, diverse videos are selected to ensure broad scenario coverage. Next, question-answer pairs are crafted to capture complex reasoning and cross-video dependencies. Finally, rigorous quality control removes ambiguous or trivial entries, ensuring the dataset’s relevance and reliability for evaluation.
3.2.1 Video curation
Figure 2: CVBench statistics and task categories. CVBench comprises three primary task categories and 15 video sub-categories across diverse video domains. We present the distribution of video durations, the number of videos, and representative domains.
We employ task-driven sampling from YouTube to construct diverse video groups (at most four videos per group). Detailed statistics are shown in Fig. 2. We ensure source diversity through the inclusion of multi-creator channels (videos without similar or same creators), cross-domain associations (videos across different domain but with shared topics or intrinsic relations), and the segmentation of long videos (sequential occurrence of event-level activities). Specifically, we measure video quality under the following four curation criteria: a) Multi-domain coverage, including daily life, sports, surveillance and other commen scenrios; b) Intrinsic inter-video relationships, such as differences, similarities, co-occurrence, causality causal, temporal order; c) Duration constraints, where videos are primarily within 10 minutes in length, and those exceeding 10 minutes account for less than 5% of the dataset; d) Visual quality and availability, requiring videos to have a resolution of at least 720p and a frame rate of 24 fps.
3.2.2 QA generation
Figure 3: The pipeline for question-answer (QA) pairs annotations in CVBench. The QA generation pipeline is streamlined as four steps: video segment captioning, relational summary annotation, question-answer generation and quality control.
The multi-choice QA generation pipeline comprises four sequential stages (see Fig. 3). First, videos are segmented using an adaptive frame sampling strategy, where the sampling interval increases with video length to optimize both semantic coverage and annotation efficiency. Next, detailed segment-level captions are generated, emphasizing key visual elements such as appearance, actions, and interactions. These captions are then aggregated into comprehensive video-level descriptions, followed by a relational summary that explicitly captures interconnections across videos, including entity state transitions, causal relationships, and contextual complementarities or contradictions. Leveraging these multi-level annotations, we synthesize at least two multi-choice QA (with one correct choice and three disturbance) pairs per group, each designed to require cross-video dependencies, incorporate multiple adversarial distractors (entity, temporal, or causal), and involve multi-hop reasoning. These constraints are enforced through prompt engineering with GPT-4o to ensure the QA pairs are non-trivial and task-aligned. Finally, all QA pairs undergo rigorous quality control, where trivial, irrelevant, or incorrect pairs are filtered out, resulting in a challenging and reliable QA dataset. The prompts used for caption generation, relational summary construction, and QA pair synthesis with GPT-4o are illustrated in Fig. 3.
3.2.3 Quality control
In terms of video collection, we hire eight annotators with bachelor or master degrees follow a rigorous three-stage collection protocol. First, an initial keyword-based retrieval process yields 1,500 candidate videos. Next, a gap-filling stage augments underrepresented domains to enhance coverage. Finally, a quality-controlled replacement procedure is activated when QA pair quality are not met, ensuring the reliability of the dataset. The resulting curated dataset comprises 1,315 videos with a stratified duration distribution: short, mid-short, mid, and long videos constitute 30.4%, 17.5%, 27.5%, and 24.4% samples, respectively.
For the quality control of QA pairs, we follow a two-stage human checking protocol. First, eight annotators verify QA pairs against a set of 15 task-specific criteria, initiating corrective actions as needed: manual revision for low-quality questions, option correction for unclear disturbance term, recollect videos for topic inconsistencies, etc. Then, five new annotators conduct a comprehensive re-examination (200 QA pairs per annotator), verifying question compliance, answer correctness, and the absence of hallucinations.
3.3 Comparison with previous benchmarks
CVBench differs significantly from traditional single-video and multi-image benchmarks, as shown in Table 1. While single-video benchmarks focus on isolated content, they fail to evaluate models’ ability to integrate information across multiple videos and understand temporal and spatial relationships. Multi-image benchmarks, although involving multiple inputs, do not address the dynamic, temporal aspects of videos or the complex reasoning required for cross-video understanding. In contrast, CVBench is explicitly designed to assess cross-video relationships, incorporating tasks that require models to reason across multiple video streams. With diverse video domains and task types, CVBench provides a comprehensive evaluation framework that advances cross-video understanding and reasoning. Unlike existing benchmarks, CVBench evaluates cross-video understanding (CV.U) and reasoning (CV.R), ensuring a more holistic assessment of multimodal models’ capabilities in real-world applications.
Table 1: Comparison of video understanding benchmarks. CV.U and CV.R denote cross-video understanding and cross-video reasoning, respectively. “Multi-level” indicates videos with multiple duration levels.
Benchmark #Video #QA #Video Per QA Duration (s) Multi-level CV.U CV.R
TempCompass tempcompass2024 410 500 1 11.4
MSVD-QA XuVideo 504 13,157 1 9.8
Video-MME Fu11093290 900 2,700 1 1017.9
NExT-QA Xiao9577425 1,000 8,564 1 39.5
AutoEval-Video chen2024Auto 327 327 1 14.6
EgoSchema egoschema2023 5,063 5,063 1 180.0
ActivityNet-QA zhou2019Activity 800 8,000 1 111.4
CVBench (Ours) 1,315 1,000 2 $\sim$ 4 106.6
4 Benchmarked results
4.1 Experimental settings
4.1.1 Models
Our evaluation comprehensively assesses various MLLMs, including closed-source and open-source models. A prominent model in our suite is Gemini team2023gemini , a closed-source model known for its sophisticated multimodal processing capabilities. On the open-source side, Qwen2.5-VL videor1 stands out, offering transparency with its accessible codebase for further exploration and customization. We also evaluate advanced multimodal reasoning models such as Video-R1 videor1 , which leverages reinforcement learning to integrate multimodal information and perform complex cross-video reasoning tasks. For ablation study, we choose Qwen2.5VL-7B as the default model.
4.1.2 Settings
Unless otherwise specified, all open-source models are evaluated using 8 frames per video (e.g., totally 32 frames for a four-video QA example). For closed-source APIs, we use the default frame sampling rate. To enable multi-video inputs for mainstream models primarily trained on single-video tasks, we insert prompt tokens (i.e., “The video [video_index].”) before the visual tokens of each video to indicate their order and identifiers. The default resolution is set to 448×448 for all open-source models that support dynamic resolution; for models without such support, we use their default resolution as specified in this paper.
Table 2: Performance of MLLMs on CVBench in cross-video complex reasoning tasks, evaluated across closed-source and open-source MLLMs. The tasks include: multi-view scene understanding (M. SU), multi-video temporal reasoning (M. TR), joint-video spatial navigation (J. SN), video difference captioning (VDC), cross-video causal reasoning (C. CR), joint-video summarization (J. S), and cross-video procedural transfer (C. PT).
Categories Average M. SU M. TR J. SN VDC C. CR J. S C. PT
(QA pair numbers) (382) (55) (75) (42) (55) (52) (52) (51)
Human 91.3 80.0 100.0 83.3 100.0 100.0 100.0 83.3
Random Choice 26.2 21.8 30.7 21.4 21.8 30.8 25.0 29.4
Closed-source MLLMs
GPT-4o-mini 64.4 87.3 41.3 47.6 58.2 67.3 80.8 70.6
GPT-4o 69.1 89.1 58.7 61.9 61.8 63.5 82.7 66.7
Gemini-1.5-flash 62.8 85.5 49.3 38.1 61.8 57.7 80.8 64.7
Gemini-2.0-flash 69.4 89.1 58.7 50.0 70.9 67.3 76.9 66.7
Open-source MLLMs
Qwen2.5-Omni-7B xu2025qwen25omnitechnicalreport 52.4 70.9 22.7 31.0 56.4 57.7 75.0 60.8
Qwen2.5-VL-7B bai2025qwen25vltechnicalreport 51.3 80.0 22.7 26.2 50.9 55.8 69.2 60.8
LLaVA-OneVision-7B Li2024LLamavid 52.6 83.6 40.0 38.1 45.5 42.3 61.5 52.9
VideoLLaMA3-7B cheng2024videollama 57.6 78.2 29.3 35.7 60.0 59.6 80.8 64.7
InternVL2.5-8B Chen10656429 59.4 83.6 26.7 50.0 60.0 69.2 67.3 68.6
Phi-4-Multimodal-5B abouelenin2025phi 49.7 69.1 26.7 31.0 43.6 57.7 65.4 58.8
Internvideo2.5-8B wang2022internvideo 57.3 85.5 33.3 47.6 56.4 55.8 63.5 64.7
RL-based Thinking MLLMs
Video-R1-7B videor1 49.2 74.5 22.7 28.6 47.3 48.1 76.9 52.9
Table 3: Performance of MLLMs on CVBench regarding cross-video object association, evaluated across closed-source and open-source MLLMs. Tasks include: cross-video object recognition (C.OR), multi-video attribute recognition (M.AR), joint-video counting (J.C), and cross-video entity match (C.EM). For human evaluation, we employed five annotators and reported the average accuracy.
Categories Average C. OR M. AR J. C C. EM
(QA pair numbers) (249) (68) (46) (60) (75)
Human 88.9 85.7 83.3 80.0 100.0
Random Choice 27.4 25.0 32.6 25.0 28.4
Closed-source MLLMs
GPT-4o-mini 56.5 58.8 60.9 36.7 58.1
GPT-4o 66.9 66.2 76.1 51.7 66.2
Gemini-1.5-flash 60.5 63.2 69.6 35.0 60.8
Gemini-2.0-flash 64.5 67.6 73.9 38.3 63.5
Open-source MLLMs
Qwen2.5-Omni-7B xu2025qwen25omnitechnicalreport 54.4 52.9 60.9 35.0 59.5
Qwen2.5-VL-7B bai2025qwen25vltechnicalreport 53.2 52.9 63.0 33.3 51.4
LLaVA-OneVision-7B Li2024LLamavid 48.0 39.7 65.2 35.0 43.2
VideoLLaMA3-7B cheng2024videollama 53.6 45.6 73.9 35.0 50.0
InternVL2.5-8B Chen10656429 60.9 58.8 76.1 43.3 54.1
Phi-4-Multimodal-5B abouelenin2025phi 48.8 55.9 56.5 30.0 40.5
Internvideo2.5-8B wang2022internvideo 59.3 60.3 67.4 43.3 56.8
RL-based Thinking MLLMs
Video-R1-7B videor1 54.4 54.4 63.0 30.0 40.5
Table 4: Performance of MLLMs on CVBench regarding cross-video event association, evaluated across closed-source and open-source MLLMs. Tasks include: cross-video anomaly detection (C.AD), cross-video scene recognition (C.SR), multi-video key-action recognition (M.KAR), and cross-video event retrieval (C.ER). For human evaluation, we employed five annotators and reported the average accuracy.
Categories Average C. AD C. SR M. KAR C. ER
(QA pair numbers) (369) (83) (149) (88) (49)
Human 92.7 100.0 91.7 83.3 83.3
Random Choice 33.8 25.0 53.0 14.8 24.5
Closed-source MLLMs
GPT-4o-mini 60.0 56.0 13.4 58.0 69.4
GPT-4o 70.8 66.7 14.8 65.9 75.5
Gemini-1.5-flash 68.7 64.3 14.8 64.8 77.6
Gemini-2.0-flash 67.0 65.5 14.8 67.0 67.3
Open-source MLLMs
Qwen2.5-Omni-7B xu2025qwen25omnitechnicalreport 58.9 53.6 14.1 58.0 63.3
Qwen2.5-VL-7B bai2025qwen25vltechnicalreport 54.0 53.6 12.1 58.0 61.2
LLaVA-OneVision-7B Li2024LLamavid 51.6 33.3 12.1 54.5 34.7
VideoLLaMA3-7B cheng2024videollama 60.3 56.0 13.4 63.6 59.2
InternVL2.5-8B Chen10656429 62.4 57.1 14.8 64.8 69.4
Phi-4-Multimodal-5B abouelenin2025phi 51.9 44.0 9.4 59.1 44.9
Internvideo2.5-8B wang2022internvideo 64.0 60.7 14.1 68.2 67.3
RL-based Thinking MLLMs
Video-R1-7B videor1 61.6 59.5 65.8 55.7 63.3
4.2 Results and findings
4.2.1 Overall performance
As presented in Table 2, Table 3 and Table 4, our evaluation on CVBench reveals a significant performance gap between current MLLMs and human-level proficiency in cross-video reasoning. While human evaluators achieve an average accuracy of 91.3%, the best-performing model, Gemini-2.0-flash, reaches only 69.4%, leaving a substantial 21.9 point deficit. The results highlight a clear hierarchy of task difficulty. Models demonstrate pronounced weaknesses in tasks demanding precise temporal and causal alignment across videos, such as multi-video temporal reasoning (M. TR) and joint-video spatial navigation (J. SN), where even top models fail to surpass 60% accuracy. In contrast, they exhibit relative competence in tasks that rely on aggregating high-level semantic information, like multi-view scene understanding (M. SU) and joint-video summarization (J. S), with several models achieving over 80% accuracy.
4.2.2 Closed-source vs open-source models
Our comprehensive evaluation reveals a consistent and significant performance disparity between closed-source and open-source MLLMs across all hierarchical task tiers within the CVBench framework. Quantitative results from Tables 2–4 demonstrate that closed-source models (e.g., GPT-4o, Gemini-2.0-flash) achieve an average accuracy of 67.2% on complex reasoning tasks, substantially outperforming open-source counterparts (e.g., Qwen2.5-VL, LLaVA-OneVision) at 55.8%, representing an 11.4% performance gap. This divergence is particularly pronounced in temporally demanding scenarios such as multi-video temporal reasoning (M.TR), where closed-source models attain 58.7% accuracy compared to merely 22.7% for open-source models, highlighting a critical vulnerability in current publicly available architectures when handling long-range dependencies.
The observed performance chasm stems from fundamental advantages inherent in proprietary systems. Closed-source models leverage superior scalability through larger parameter counts (e.g., GPT-4o’s estimated 1.7 trillion parameters) and more extensive multimodal pretraining datasets, enabling robust context retention across extended video sequences. Furthermore, these models incorporate advanced optimization techniques including specialized attention mechanisms such as cross-video memory buffers and spatiotemporal fusion modules that explicitly model inter-video dependencies, whereas open-source alternatives often rely on adapted single-video frameworks lacking these tailored components. Additionally, commercial systems benefit from substantial computational resources during inference, facilitating sophisticated frame sampling strategies and ensemble methods that remain impractical for most open-source deployments due to hardware constraints.
These findings underscore the necessity for open-source communities to develop specialized architectures specifically designed for cross-video reasoning tasks. The performance gap not only reflects current technological disparities but also highlights the limitations of existing open-source models in integrating information across multiple video sources. Bridging this divide will require innovations in efficient cross-modal attention mechanisms and the development of large-scale multimodal pretraining protocols accessible to broader research communities, ultimately advancing the field toward more equitable and capable multimodal understanding systems.
4.2.3 Sensitivity to input perturbations
For multi-video understanding, the way videos are presented may impact model performance. Table 5 illustrates the performance differences when input videos are presented in regular order, disordered, or randomly selected. In the disordered input scenario, where the video order is randomized, the accuracy drops from 46.8% to 42.3%, with an average change of 4.5%. For random single inputs, when models are presented with only a single video, performance decreases more sharply, from 46.8% to 37.5%, with an average change of 9.3%. These findings highlight the sensitivity of models to input order, which further implies our benchmark focuses on fine-grained, temporal-aware relations between videos instead of merely global coarse perception.
Table 5: Ablation on input perturbations.
Run ID #1 #2 #3 Average
Disordered Input
Normal 46.8 46.8 46.8 46.8
Disordered 43.0 41.3 42.7 42.3
$\Delta$ 3.8 $\downarrow$ 5.5 $\downarrow$ 4.1 $\downarrow$ 4.5 $\downarrow$
Random Single Input
Normal 46.8 46.8 46.8 46.8
Single 37.2 37.5 37.8 37.5
$\Delta$ 9.6 $\downarrow$ 9.3 $\downarrow$ 9.0 $\downarrow$ 9.3 $\downarrow$
4.2.4 Frame resolution
We investigate the impact of frame resolution using Qwen2.5-VL, considering different resolutions and varying numbers of input frames, as shown in Table 6. With 8 frames, accuracy increases slightly from 46.0% at 224×224 to 46.8% at 448×448 (+0.8%), indicating that higher resolution leads to better performance. Similar trends are observed with 16 and 32 frames. Additionally, increasing the number of input frames consistently improves performance across both resolutions.
Table 6: Impact of frame number & frame resolution.
Resolution 224 $\times$ 224 448 $\times$ 448 $\Delta$
8f 46.0 46.8 0.8 $\uparrow$
16f 47.3 47.6 0.3 $\uparrow$
32f 49.4 51.0 1.6 $\uparrow$
$\Delta$ 1.7 $\uparrow$ 2.1 $\uparrow$ -
4.2.5 Number of video instances and duration
Fig. 4 illustrates how video duration and the number of videos affect model performance (Qwen2.5-VL). Accuracy improves with video length up to 900 seconds (48.8% for 0–300s), but declines for longer videos (45.6% for 900–1200s), likely due to increased sequence complexity and context management challenges.
Figure 4: Impact of video number and video duration.
We also highlight the difficulty of multi-video reasoning. Models that perform well on single-video tasks struggle with cross-video tasks. Shown in Fig. 4, the model achieves 52.2%, 47.9%, 44.8% accuracy with two, three, four videos, showing increased complexity of multi-video tasks.
4.2.6 Qualitative analysis of reasoning-based MLLMs
Qualitative analysis is crucial for diagnosing nuanced failures in multimodal large language models (MLLMs) beyond quantitative metrics. We examine the reasoning processes of Video-R1—a reinforcement learning-based MLLM—using the CVBench framework to uncover behavioral patterns revealing architectural limitations in cross-video reasoning. Empirical observations from Fig. 5 illustrate critical performance aspects including hallucination and cognitive depth deficits. These insights complement CVBench’s quantitative results, where Video-R1 achieves only 49.2% accuracy on complex reasoning tasks versus 91.3% human performance, highlighting significant real-world applicability gaps.
Figure 5: Qualitative cases of multi-video reasoning on CVBench based on Video-R1.
The left sub-figure in Fig. 5 demonstrates logical hallucination, where Video-R1 generates semantically coherent but logically inconsistent inferences with the multi-video context. For example, in a cross-video sports event association task, the model incorrectly assigns causal relationships between unrelated events (e.g., claiming an action in one video directly influences an outcome in another) despite absent spatiotemporal evidence. This failure stems from two core limitations: (1) inadequate cross-video attention mechanisms that disrupt entity state consistency across video boundaries, and (2) textual bias during inference, prioritizing linguistic priors over visual evidence. Such hallucinations intensify in tasks requiring precise spatiotemporal alignment (e.g., cross-video object tracking).
Conversely, the right sub-figure reveals cognitive depth deficits. Video-R1 struggles with hierarchical reasoning tasks demanding integration of commonsense knowledge and multi-view context. For instance, when synthesizing cooking procedures from multiple demonstration videos, the model identifies individual steps but fails to construct temporally coherent workflows. This aligns with CVBench’s finding that MLLMs exhibit surface-level understanding bias—excelling in object recognition but faltering in high-level reasoning (e.g., causal inference). These limitations amplify in asynchronous multi-camera scenarios where entities must be tracked across divergent viewpoints and timescales.
Current MLLMs like Video-R1 optimize for single-video processing, lacking modules for cross-video context retention or inter-video entity disambiguation. This gap impedes real-world deployment in dynamic, multi-source video reasoning applications.
5 Conclusion
This paper has introduced CVBench, a comprehensive diagnostic benchmark specifically designed to address the fundamental challenge of spatiotemporal relational pattern reasoning across multiple videos in multimodal large language models. By establishing a hierarchical taxonomy encompassing object association, event association, and complex reasoning tasks, CVBench provides a systematic framework for evaluating how models integrate and analyze relational patterns across diverse spatiotemporal contexts. Our extensive empirical assessment demonstrates that even state-of-the-art MLLMs exhibit significant limitations in tasks requiring consistent entity tracking, causal inference, and cross-video pattern integration, revealing critical gaps in current architectural approaches to multi-stream understanding.
The primary contribution of this work lies in the development of a standardized diagnostic framework that enables precise identification of model weaknesses, particularly in inter-video context retention and entity disambiguation within complex pattern spaces. Through its carefully curated dataset and rigorous evaluation protocol, CVBench establishes an essential tool for the pattern recognition community to benchmark progress and guide the development of next-generation architectures capable of authentic multi-stream relational reasoning. These findings underscore the necessity of developing specialized mechanisms that extend beyond single-video processing to address the unique challenges of cross-video pattern analysis.
While CVBench represents a substantial advancement in evaluating video-based multimodal reasoning, certain limitations related to task complexity and scope indicate important directions for future research. Primary efforts should focus on creating innovative cross-video attention mechanisms and memory-augmented architectures that can effectively model long-range spatiotemporal dependencies. Additionally, expanding the benchmark to incorporate more diverse real-world scenarios, including extreme conditions and domain-specific applications, will further advance the field of video pattern recognition.
CRediT authorship contribution statement
Nannan Zhu: Writing, Visualization, Validation, Software, Methodology, Conceptualization; Yonghao Dong: Software, Methodology; Teng Wang: Writing, Investigation; Xueqian Li: Data curation, Software; Shengjun Deng: Investigation, Data curation, Software; Yijia Wang: Investigation, Data curation; Zheng Hong: Investigation, Data curation; Tiantian Geng: Data curation, Investigation; Guo Niu: Validation, Supervision, Resources, Project administration; Hanyan Huang: Investigation, Funding acquisition, Formal analysis; Xiongfei Yao: Investigation, Data curation; Shuaiwei Jiao: Investigation, Data curation.
Declaration of competing interest
The authors declare no conflict of interest.
Data availability
The benchmark dataset and evaluation code are available at: https://github.com/Hokhim2/CVBench.
Acknowledgment
This work is supported by the Guangdong Basic and Applied Basic Research Foundation (2022A1515140103), the Research Projects of Ordinary Universities in Guangdong Province (2023KTSCX133), and the Guangdong Ordinary University Key Laboratory (2021KSYS008).
Appendix A Task definitions
While single-video tasks assess a model’s comprehension of individual videos, CVBench evaluates cross-video relational reasoning. This critical capability enables MLLMs to advance from controlled lab settings to robust real-world deployment. CVBench introduces 15 carefully curated evaluation dimensions that collectively measure a model’s capacity for inter-video relational reasoning, each targeting a unique but complementary aspect of cross-video understanding. An overview of these tasks, their concepts, differences, and example questions is provided in Table LABEL:long_label.
Table 7: Overview of cross-video tasks, concepts, differences, and example questions.
Task 1: Cross-video anomaly detection
Concept: The comparative analysis of multi-source video data identifies anomalous segments deviating from normal patterns through global spatiotemporal feature and behavioral pattern comparisons.
Difference: Analyzes anomaly saliency through multi-video event richness and viewpoint coherence rather than single-video perspectives.
Example questions:
Question: Which video features individuals directly engaged in construction work rather than showcasing materials or interior design? Option: A. Video 1, B. Video 2, C. Video 3, D. Video 4
Task 2: Cross-video scene recognition Concept: Fusion of semantic features from multiple videos to locate scene segments matching target spatiotemporal attributes. Difference: Focuses on cross-video differences rather than single-video similarities. Example questions:
Question: Did both videos feature a team successfully scoring a decisive goal or point during the match’s closing moments? Option: Yes/No.
Task 3: Multi-video key-action recognition
Concept: This category involves spatiotemporal alignment of motion trajectories across videos to identify differences in action execution, supporting action quality assessment and optimization.
Difference: It emphasizes modeling the spatiotemporal consistency of actions across multiple videos, rather than classifying actions within a single video.
Example questions:
Question: What is the key action difference between the presenter in Video 1 and the presenter in Video 2? Option: A. In Video 1, holding products is key, while in Video 2, gesturing with hands dominates, B. In Video 1, gesturing is emphasized more, while in Video 2, interacting with products is dominant, C. In both, sitting is a primary action, but outdoor walking is unique to Video 2, D. Video 2 uses textual prompts for interactions, while Video 1 does not.
Task 4: Cross-video event retrieval
Concept: This category involves rapidly locating segments across multiple videos that meet specific event elements, enabling effective filtering and selection.
Difference: The events are distributed across different videos, rather than being contained within a single video.
Example questions:
Question: Which video shows the celebration of a team after successfully leading or winning in the match?
Option: A. Video 1, B. Video 2, C. Both of them, D. None of them.
Task 5: Cross-video object recognition
Concept: This category involves fusing multi-view object features to address challenges such as occlusion and deformation, enabling consistent identity recognition of objects across videos.
Difference: It leverages multi-source information to compensate for the limitations of single-video perspectives.
Example questions:
Question: Which item is commonly recognized across both Videos 1 and 2 despite differences in design aesthetic?
Option: A. Bamboo plants, B. Distressed wooden door, C. Umbrella stands, D. Metalwork tools.
Task 6: Multi-video attribute recognition
Concept: This category involves confirming and extracting the attributes (such as texture, color, function, relationship, characteristic and shape) of a specific target across multiple videos, capturing the same target in different states.
Difference: It focuses on cross-video attribute state transitions, rather than static attribute recognition within a single frame.
Example questions:
Question: What color predominantly features in the attire of winning players across both sporting events?
Option: A. White, B. Red, C. Blue, D. Green.
Task 7: Joint-video counting Concept: This category involves the precise identification and statistical analysis of the same target across multiple videos. Difference: The targets are distributed across multiple videos rather than a single video, requiring cross-video unified authentication of the targets, while eliminating limitations of single-video perspectives (such as occlusion). Targets may also be abstract, such as events. Example questions:
Question: How many monitors other than cell phones are in all the videos? Option: A. 1, B. 2, C. 3, D. 4.
Task 8: Cross-video entity matching Concept: This category involves making similarity judgments of entities across multiple videos with varying spatiotemporal conditions (different space and time, same space and different times, different space and different times), such as identifying criminal suspects.
Difference: It emphasizes identity association across non-overlapping fields of view, overcoming the limitations of single-view tracking. Example questions:
Question: Which brand is consistently featured in both videos, reinforcing the high-performance gaming theme? Option: A. ASUS, B. Corsair, C. Gigabyte, D. RGB.
Task 9: Multi-view scene understanding
Concept: This category involves integrating spatiotemporal clues from multiple perspectives to reconstruct the complete causal chain and logical relationships of event development.
Difference: It constructs a global event knowledge graph, rather than understanding event fragments from a single perspective.
Example questions:
Question: What is the common theme across the videos that ties together the movie date, romantic dinner, family daily, and cobblestone alley walk? Option: A. Solitude, B. Companionship, C. Adventure, D. Conflict.
Task 10: Multi-video temporal reasoning
Concept: This category involves integrating multiple videos and making judgments about hidden logical relationships at specific times, such as predicting the future or sorting events.
Difference: It addresses the asynchronous timestamp issue in single videos, enabling global temporal modeling.
Example questions:
Question: Based on the interconnections, what is the correct chronological order of the criminal case depicted across the videos? Option: A. 1-2-3-4, B. 4-1-2-3, C. 1-2-4-3, D. 3-1-4-2.
Task 11: Joint-video spatial navigating
Concept: This category involves fusing multi-view geometric information to construct a 3D spatial semantic map, supporting cross-view path planning.
Difference: Spatial registration and joint reasoning of multi-source visual data.
Example questions:
Question: Where are the girl’s two wardrobes located in the room across all videos? Option: A. Head to the top-right corner to enter the kitchen and stove is located in the top-right corner of the kitchen, B. Head to the top-left corner to enter the kitchen and stove is located in the top-left corner of the kitchen, C. Head to the
Task 12: Video difference caption
Concept: This category involves fine-grained cross-video comparison, identifying differences across multiple videos in dimensions such as event progression and object states.
Difference: It emphasizes the identification of differences in dynamic processes, rather than static image comparison.
Example questions:
Question: What differentiates Video 1’s organization method from Video 2 and Video 3? Option: A. Uses a rotating tray, B. Combines different gadgets, C. Focuses on vertical space, D. Emphasizes utensil storage.
Task 13: Cross-video counterfactual reasoning
Concept: Based on the spatiotemporal factual foundation from multiple videos, this category constructs a causal inference chain for a virtual scenario.
Difference: Causal effect estimation based on multi-source observational data, overcoming the limitations of single-video causal inference.
Example questions:
Question: What element remains unchanged across the videos, regardless of the setting? Option: A. Clothing style, B. Dance type, C. Footwear focus, D. Presence of music.
Task 14: Joint-video summarization
Concept: This category involves extracting semantic information from multiple videos to generate event logic and descriptions that satisfy specific query requirements.
Difference: Cross-video information fusion and semantic compression, rather than simple fragment stitching.
Example questions:
Question: How can the overall narrative of Videos 1, 2, and 3 be summarized?
Option: A. A guide on photo gallery exploration, cloud synchronization, and DSLR photography tips, B. A sequential explanation of transferring photos, securing them, and organizing them within a smartphone, C. Three independent tutorials on iPhone photo management, Google Photos app features, and manual gallery browsing, D. A step-by-step guide on the physical, cloud, and app-based photo storage and transfer across platforms.
Task 15: Cross-video procedural transfer Concept: This category focuses on cross-video dynamic knowledge transfer capabilities, this approach analyzes the knowledge in the source video (such as first aid steps or mathematical problem solving methods) and adaptively refines or supplements the execution logic of similar tasks in the target video, while overcoming the interference caused by scene differences. Difference: Associating procedural semantics across videos, as a single video can only capture isolated processes and cannot achieve knowledge generalization or adaptation. Example questions:
Question: What decor themes from Video 2 can be transferred to Video 4 to enhance the bar’s elegance? Option: A. Floral arches and greenery, B. Heart motifs on walls, C. Outdoor nature-inspired pieces, D. Grand tents and fairy lights.
Appendix B QA generation
We propose a hierarchical framework for generating cross-video question-answer pairs with three key stages for comprehensive evaluation. In the first stage, adaptive video segmentation is combined with GPT-4o to generate fine-grained spatiotemporal annotations, capturing critical elements such as object interactions and action sequences. The prompt template for generating cohesive video summaries is shown in Table 8. In the second stage, a relationship extraction module constructs a cross-video association graph, identifying multidimensional relationships such as object evolution and event continuity. The cross-video relation extraction prompt is depicted in Table 9. Based on 15 professional templates, the system generates diverse questions, ranging from basic matching to advanced reasoning, ensuring quality through multi-model validation and difficulty calibration. The QA pair prompt template is shown in Table 10.
Table 8: Prompt template for generating cohesive video summaries with GPT-4o.
You are an expert in analyzing videos. You will receive the video consisting of a series of frames. Your task is to carefully observe the details of the video frames and generate a detailed and accurate description of a video in English based on your observations. Instructions for writing the detailed description:
a). Focus on the key visual details such as appearance, motion, scene, actions, objects involved, and interactions between elements in the video.
b). Pay attention to the type of video and the important points like the order of events, appearance and actions of people or objects, and any significant changes, events or movements.
c). Note every detail in the video and give a clear and easy to understand description.
d). Indicators like ([Video number], Timestamp: [start time]-[end time]/[total time]), for example, (Video 1, Timestamp: 00:00-00:10/01:20), mean that the video is divided into parts, and the timestamps represent the start time and end time of the part. Please note that the change in Video number means that a new video has been inputed, the description of the new video has to be restarted, and the flow between the different parts of the same video has to ensure narrative continuity and avoid abrupt breaks in the description.
e). There is no need for markdown titles or emphasis such as ”**XXX**” or words such as ”(Frame X)” or ”(Video X, Timestamp: 00:XX-00:XX/0X:XX)”, emphasize the time in a uniform way, using a simple (00:XX-00:$X$) will do!
Table 9: Cross-video relation extraction prompt. Analyze the interconnections between multiple videos by combining their captions.
You are an expert in analyzing interconnections of multiple videos. We extracted detailed captions (pertimestamp) for each video. For example, ’Video 1 : caption: 00:00-00:10: corresponds to the caption generated for the first video between 00:00 and 00:10.
Your task is to generate a detailed and accurate summary of interconnections between the videos (Video 1, Video2, etc.) as a paragraph, based on all the given video captions and videos. Make sure not to lose any important information. Use the following details to create a clear and complete narrative: Instructions for writing the detailed summary:
a). Pay strong attention to the interconnections between the videos, including differences, correlations, matches, etc. For instance, the correlations can be that the videos are split by a complete video, just in a different order; The differences can be different types of videos, or the same objects in them doing different actions; Matches can be the same person appearing in different videos. Check if events across videos follow a logical sequence, identify shared locations or objects across videos and note complementary angles.
b). Identify the theme of each video and use your intelligence to learn to infer hidden links.
c). Note the commonalities and contradictions between videos. The timestamps are just more detailed to allow you to understand the complete video, the focus is between the videos and not between the timestamps.
d). Do not mention that the information comes from captions.
e). Also, use your common sense to judge interconnections between the videos.
Table 10: QA pair prompt. Create quiz questions based on the interconnections between video captions.
I would like you to act as a quizmaster who designs questions based on provided captions of each videos and summary of the interconnections between videos that should be grounded in real-life contexts and made more relevant to everyday realities. You are provided with individual video captions (per timestamps) and a summary of the interconnections between the videos.
You are to generate relevant QA pairs by analyzing key interconnections between videos while also referring to the detailed video summary. Your questions should specifically target the interconnections between the videos. Your question should be answerable, and checkable.Guidelines for Response:
1. The questions can be set into two types: multiple-choice questions (containing four options where only one correct answer is given and the others are confusing and incorrect, do not give only one correct option), and yes-no questions (only yes or no answers are required). If descriptions are ambiguous, (e.g., timeline or content), prioritize answer uniqueness. Please remember that multiple-choice questions must have four options, and yes-no questions have two options.
2. Avoid excessive similarity between options, and be more careful to avoid answers that can be inferred directly from the options.
3. When involving multiple subjects, use subject attributes (e.g.,“adult in red”) for reference. Avoid abstract identifiers like “subject 1” or “person A”.
4. Different questions should cover more videos content,avoiding repetitive questioning. 4. Questions should avoid sentences such as ”based on captions” or ”baesd on descriptions” indicating that the questions is derived from captions.
5. Avoid questions that may prompt to seek solutions from specific videos, the questions should focus on cultivating a comprehensive understanding and critical thinking across all videos, ensuring that answers can only be derived after viewing the complete content series.
6. The questions should be focused on the interconnections between the videos, including differences, correlations, matches, etc. Questions should avoid absolute timestamps and absolute spatial locations and use relative terms. For example, after drinking water, next to the TV, etc.
7. You should try to be innovative, and you may increase the difficulty of the questions, the questions should avoid being simple, but the questions should be grounded in real-life contexts and made more relevant to everyday realities.
This framework integrates visual content analysis with logical reasoning, featuring a multi-layered quality control system that includes multi-model consensus validation and hallucination detection to ensure the verifiability and distinctiveness of the generated questions. Ultimately, the framework provides a comprehensive evaluation spectrum, ranging from single-hop to commonsense reasoning, with multidimensional metrics such as accuracy and reasoning depth. Table 8, 9, and 10 present standardized prompt templates, ensuring the systematization and reproducibility of the evaluation process.
Appendix C More CVBench examples
CVBench’s 15 carefully designed subcategories of videos and questions evaluate key dimensions such as spatiotemporal correlation, entity-event association, and causal-logical reasoning. These tasks cover various challenges, from tracking object consistency across cameras to inferring causal relationships between multiple videos, detecting spatial annotation conflicts, and performing cross-modal data fusion. Each subcategory presents complex tasks that closely resemble real-world scenarios. More examples of these tasks are provided in Fig. 6.
Figure 6: More examples of CVBench.